You’ve observed the following returns on Crash-n-Burn Computer’s stock over the past five years: 4 percent, -10 percent, 28 percent, 21 percent, and 12 percent.

  

Requirement 1:
What was the arithmetic average return on Crash-n-Burn’s stock over this five year period?
(Click to select)  8.91%  11.99%  11.00%  13.09%  13.75%

   

Requirement 2:
(a)

What was the variance of Crash-n-Burn’s returns over this period? (Do not round intermediate calculations.)

(Click to select)  0.02200  0.01782  0.02295  0.02002  0.01760

   

(b)

What was the standard deviation of Crash-n-Burn’s returns over this period? (Do not round intermediate calculations.)

(Click to select)  12.01%  15.15%  17.65%  14.83%  13.27%
Answer #1

arithmetic average return=Total return/Total time period

=(4-10+28+21+12)/5

=11%

Return (Return-Average Return)^2
4 (4-11)^2=49
-10 (-10-11)^2=441
28 (28-11)^2=289
21 (21-11)^2=100
12 (12-11)^2=1
Total=880%

Standard deviation=[Total (Return-Average Return)^2/(Time period-1)]^(1/2)

=[880/(5-1)]^(1/2)

=[880/4]^(1/2)

=14.83%(Approx).

Variance=Standard deviation^2

=0.02200